A researcher orders a broth of 8.5% glucose for her lab. However, she needs a stronger broth, one that is 25.9% glucose. Fortunately, she has 26.9 liters of 35.1% glucose broth in the stock room. How much 24.9% glucose broth can she make?
Answer by ankor@dixie-net.com(12701)

A researcher orders a broth of 8.5% glucose for her lab. However, she needs a stronger broth, one that is 25.9% glucose. Fortunately, she has 26.9 liters of 35.1% glucose broth in the stock room. How much 24.9% glucose broth can she make?
:
You write the resulting amt as 25.9% then you write it as 24.9%, I am assuming
that 25.9% is the correct one
:
Let x = amt (liters) of 8.5% broth required
:
The total resulting amt = (x+26.9) liters
:
Put all percents into decimal form
:
.351(26.9) + .085x = .259(x+26.9)
:
9.4419 + .085x = .259x + 6.9671
:
9.4419 - 6.9671 = .259x - .085x
:
2.4748 = .174x
:
x = 2.4748/.174
:
x = 14.223 liters of 8.5% stuff
Therefore:
14.223 + 26.9 = 41.123 liters of 25.9% can be made
:
:
Check solution;
:
.351(26.9) + .085(14.223) = .259(41.123)
9.4419 + 1.209 = 10.65
:
Did this help? If I assumed the wrong percentage, you can use this method, just change 25.9% to 24.9%